### v2

@@ 33             $A_{ij}^{Trim}=1$ ;
Algorithm 1 Causally graph Attention network for Trimming heterophilic graph(CAT)
@@ 5.2 Class-level Semantic Clustering Module
Based on the Class-level Semantic Space Hypothesis, we consider that for the central node, the neighbors impact the self-attention learning with their class as the basic unit. It is very intuitive on heterophilic graphs, when the representations of graph nodes are hard to distinguish, the more nodes we observe for a class, the more beneficial it is to get the global distribution of that class.
In that semantic space, we treat the local neighbors of same cluster as a whole, which is referred to as Semantic Cluster $SC=f_{\text{clustering }}(x),\quad x\in\mathbb{R}^{n\times F},\quad SC\in\mathbb{R}^{n\times 1}$. Where $SC(i)\in C$ represents the cluster class of nodes with the index $i$. Accordantly, the center of each SC in that semantic space is $SC_{c}=\left\{\frac{\sum\limits_{SC(v)=c}SC_{v_{1}}}{n_{c}},\frac{\sum\limits_{SC(v)=c}SC_{v_{2}}}{n_{c}},\ldots\frac{\sum\limits_{SC(v)=c}SC_{v_{d}}}{n_{c}}\right\},c\in C$. We can update the causal graph proposed in Figure 2 to Figure 9.
Figure 9: Causal graph behind GAT at the level of Semantic Cluster.
As shown in Figure 10, three learning paradigms can be adopted in this module to obtain Class-level Semantic Clusters, ordered from low to high prior knowledge of the node category distribution they are: unsupervised, semi-supervised, and supervised learning. The more information we acquire about the categorical distribution of graph node, the closer the Semantic Cluster distribution we obtained will be to the distribution in the ideal semantic space. This further indicates a more accurate estimation of the Total Effect of attention learning on Class-level Semantic Cluster corresponding to each category. We construct three CAT variants by adopting following three learning paradigms in this module:
  • Unsupervised manner: For all nodes in the graph, their categorical labels are unseen. Unsupervised clustering methods can be employed to obtain a rough semantic space with the input of node features. The CAT variant under this manner is referred to as CAT-unsup.
  • Semi-supervised manner: Categorical labels for a fixed ratio of nodes are known and used to inference the labels of unknown nodes. Classification methods with a semi-supervised setting can be employed to obtain a fewer rough semantic space with the input of node features. The CAT variant under this manner is referred to as CAT-semi.
  • Supervised manner: Categorical labels for all nodes are known and their categorical distribution is completely and accurately observed. It should be noted that the label information is only available in the Class-level Semantic Clustering stage and will not be used for node classification. The CAT variant under this manner is referred to as CAT-sup.
@@ 5.3 Total Effect Estimation Module
As illustrated in Figure 9, there are two paths from class-level Semantic Clusters to the central node’s attention distribution that will jointly influence the representation learning of the central node. Therefore, we employed Total Effect as a measurement of Distraction Effect based on the preliminary of Causal Inference. This module contains three important steps, Semantic Cluster Intervention, Intervened Attention Leaning and Graph Trimming.
@@
    Semantic Cluster Intervention. As detailed in Section 3, Total Effect is estimated based on the intervention.This step is theoretically equivalent to asking the central node to think about a causal question: how will my attention distribution change if Semantic Cluster $c$ is removed from my LND? From the physical intuition behind this intervention-related question, it equals to a operation that renders the nodes belonging to Semantic Cluster $c$ invisible to the central node. As illustrated in Figure 11, it can be mathematically modeled as Eq.7, where $A$ represents the adjacency matrix of original graph, $\hat{A}$ represents that of intervened graph.
@@
    Intervened Attention Learning. Since the intervention on Semantic Cluster does not affect the shape of Class-level Semantic Space (which is totally governed by the data generation mechanism of the graph), we need to guarantee that before and after the intervention, the GAT model only changes the attention distribution, while the other capabilities remain unchanged. From the perspective of model implementation, we do not alter the parameters responsible for transforming node features, allowing the model to solely relearn the allocation of attention to neighboring nodes and itself. The attention assigned to Semantic Clusters can be represented in Eq.8, where $\alpha_{SC}^{c}=\sum\limits_{SC(V_{j})=c}\alpha_{ij}$, and the self-attention of central node $V_{i}$ is $\alpha_{self\_attention}=1-\sum\limits_{C}\alpha_{SC}^{c}$.
@@
    Graph Trimming. According to the concept in Section 3, we can calculate the TE on Central node’s self-attention of Semantic Cluster $c$ according to Eq.9.
@@
    The lower the value of the sum of $TE_{\alpha_{SC}}$, which implies the higher value of $TE_{\alpha_{self\_attention}}$, the more it can distract the central node and lead to the low self-attention of the central node, and vice versa. Therefore, we remove the Semantic Cluster with lower $TE_{\alpha_{SC}}$ and only retain the Semantic Cluster with the highest $TE_{\alpha_{SC}}$. In other words, only the Semantic Cluster with the lowest TE on self-attention of the central node will remain. Eventually, we obtain the adjacency matrix of the Trim Graph denoted as Eq.10, which equals to a operation that remove the adjacent edges between Distraction Neighbors and central nodes.
@@ 6.1 Databases
To ensure the richness and representatives of the data, we selected seven heterophilic graphs of three different sizes, each with an Edge Homophily lower than 0.23. The basic information of the dataset is given in the Table 2. The Edge Homophily is defined as Eq.11, and lower Edge Homophily indicates higher heterophily of graph.
@@
  • Small-size datasets. We use the WebKB dataset [42] constructed from the WebKB webpage [58]. They are collected from computer science departments of Cornell, Texas and Wisconsin university. These datasets are built from the hyperlinks between webpages, and the features of node are the bag-of-words representation. Nodes belong to five categories.
  • Medium-size datasets. We use the Chameleon and Squirrel [59] dataset collected from Wikipedia, they are applicable for node regression and node classification [42] tasks. In these datasets, the nodes represent webpages and the edges are links between them. When applied for node classification task, the target is to predict 5 classes based on the average traffic of webpages.
  • Large-size datasets. We use the Roman-empire [60] constructed from Wikipedia article. In this dataset, the nodes represent words in the text and the edges are constructed from their context. The target is to predict 18 classes based on the syntactic role of the nodes.
@@ 6.2 Experiments
We aim to explore the effect of neighboring nodes on the central node’s attention learning within the GAT mechanism. Therefore, we take the GAT with fixed architecture (which can be regarded as having the same aggregation and feature transformation ability) as the base model, and compare the discrimination ability on graph with different LND. We focus more on the difference caused by attention distribution learned by GATs, thus choosing not to tune the architecture of GATs through careful optimization and parameter tuning. In this paper, we adopted three GATs as the base model and constructed corresponding CAT variants. To distinguish the results on different base models, we replace the ’G’ in the base model name with ’C’ to represent the corresponding CAT. The base GATs and their fixed architecture are as followed:
  • GAT[25]. The originally proposed Graph Attention Network. We set the layer of GAT to 2 and multi-head number to 8.
  • GATv2[61]. A dynamic graph attention variant that can learn dynamic attention by simply switching the order of internal operations in GAT. We set the layer of GATv2 to 2 and multi-head number to 8.
  • GATv3[47]. A new attention mechanism that calculates query and key with other GNN models in GAT framework. We set the layer of GATv3 to 2 and adopt one-layer GCN in K and Q module. To better investigate the effect of attention learning, we fix the weight of calculated attention to 1 and abandon the original weighted attention strategy.
For all base GATs and their CAT variants, we use the Adam optimizer with a learning rate of 0.001 and a weight decay of 0.0001 to train the model. A single Nvidia 2080Ti is used for training with a negative log likelihood loss. The maximum iteration is 600, and the tolerance in the early stopping strategy based on the classification accuracy on the validation set is set to 50. For the evaluation of the model accuracy, we divide the dataset into training, validation and test set with the ratio of 6:2:2, and use the average classification accuracy and standard deviation of the test set in 100 repetitions as the final evaluation metrics. We set the dimensions of the hidden layers to {16,32,64,128} and adopt the optimal classification accuracy. We conduct comparison experiment and ablation experiment to verify the validity of the architecture and single modules of CAT, respectively. Visualization experiments are also carried out for further interpretation of the results.
@@ 6.2.1 Comparison Experiment
We feed the original heterophilic graph and the trimmed graph obtained by CAT variants to the base GAT to obtain the final node classification accuracy, respectively. Trimmed graphs are obtained in three variants of CAT with the following setting:
@@
  • CAT-unsup. Since the number of Semantic Clusters is known (equal to the number of target classes), we use the K-means++ algorithm in the Class-level Semantic Clustering Module for the unsupervised manner. To avoid the influence of the initial clustering centers on the results, we use 0,10,100 as random seeds respectively for the initialed clustering centers in K-means++.
@@
  • CAT-semi. For the semi-supervised manner, we employed a two-layer Multi-Layer Perception (MLP) to learn the categorical distribution of nodes. To keep the consistency of the semi-supervised node classification task, we use the same dataset split described in Section 6.1 for the MLP.
@@
  • CAT-sup. For the supervised manner, we directly use the labels to generate the Class-level Semantic Clusters.
@@
The results are shown in the Table 3. It can be found that our approach demonstrates improvements across all base GAT models. Even on the Large-size dataset with the Edge Homophily of only 0.04, the minimum of relative improvements for GAT and GATv2 achieve 13.5%, 10.1%, respectively. Adopting semi-supervised and fully-supervised paradigm can further enhance the improvement. GATv3 exhibits the best performance among all base models due to its incorporation of a new attention mechanism that calculates query and key from other GNN models, enhancing its discrimination capability. Despite this, CAT framework can further improve its classification accuracy. Among all base models, the relative improvement of CAT on GATv3 is the lowest. The reason is GATv3 already boasts a comparatively high level of discrimination ability and classification accuracy on heterophilic graphs, making further enhancement more challenging.
@@
In terms of the standard deviation of the prediction accuracy, on Small-size datasets, the deviation of CATs is relatively large as the base GATs, which is caused by the smaller size of the dataset. However, on Medium-size and large-size datasets, CAT is able to significantly reduce the deviation and achieve more stable and statistically significant predictions on the dataset.
@@
For all base GATs, CAT-sup can generally achieve the best performance than CAT-unsup and CAT-semi. It is because the more information is leveraged in the Class-level Semantic Clustering Module there by a more accurate distribution of Semantic Clusters is obtained. In the same way, this speculation can also explain why CAT-unsup performs the worst, and the CAT-semi consistently performs at a middle level. On one hand, this indicates that precise Class-level Semantic Clustering can assist the central node in better allocating attention. On the other hand, it poses a challenge to learning a more optimal Semantic Space. CAT-unsup with different random seeds also obtain significantly different performance. For example, although all CAT-unsup can outperform GAT on the Wisconsin dataset, CAT-unsup with random seed value of 10 gains over 10% lower accuracy than that with value of 100. CATv2-unsup exhibits a similar pattern on the Texas dataset. This indicates that we can barely guarantee that the learned Class-level Semantic Space is optimal or approaching optimal for the sake of unsupervised manner. Also, it indicates that the output of Class-level Semantic Clustering has a significant effect on the whole method.
@@
Table 3: Node classification accuracy. 0,10,100 represent the corresponding random seed used in the unsupervised clustering method. The best result and worst result achieved by CAT variants are marked in bold and wave line separately, and their relative improvements compared to GAT are shown below. The base models and the worst result among all models are marked in red. OOM represents out of memory.
@@ 6.2.2 Ablation Experiment
To investigate the effectiveness of components in proposed method, we conducted ablation studies on two modules separately and obtain two corresponding trimmed graphs as input. For the sake of better comparison, we selected CAT-unsup in this section because it performs the worst among three variants of CAT thereby can better support the ablation experiments. The result of ablation experiment is shown in Table 4.
@@
    CAT (random_cluster). To investigate the effectiveness of the Class-level Semantic Cluster module, we replace it with a random assigned cluster module. We set the random seed of 0, 10, 100 respectively for the random cluster assignment.
@@
    CAT(high_distraction). To investigate the effectiveness of the Total Effect Estimation, we remove the neighbors with lower distraction and create a High Distraction and Low Self-attention scenario. This model reserves the $\max\left(TE_{\alpha_{self\_attention}}\right)$ in the Total Effect Estimation.
Table 4: The ablation experiment result. The CAT in this table represents the CAT-unsup. The best accuracy is marked in bold, and the worst is in wavy line.
@@
It can be found that CAT always achieves the best performance, while CAT (high_distraction) gets the worst performance. This comparison supports our Low distraction and High self-attention assumption, and also proves the effectiveness of our Total Effect Estimation Module. CAT (random_cluster) gets the medium performance, indicating the significance of our Class-level Semantic Clustering Module, and to some extent it can help us learn the effect of each class on the performance of the central node in the semantic space. In addition, it also shows that Total Effect Estimation Module has a larger and more stable contribution to CAT.
However, we also notice that CAT (random_cluster) is able to achieve results comparable to or even exceeding CAT in very few cases, which suggests that the clustering results obtained by the Class-level Semantic Clustering Module need to be optimized, whereas random cluster may be better. This phenomenon is more striking on small-size dataset, and the possible reason is that this kind of dataset itself has class imbalance (as shown in Figure 12), which will raise the difficulty of clustering.

### v3

@@ 33             $A_{ij}^{Trim}=1$ ;
Algorithm 1 Causal graph Attention network for Trimming heterophilic graphs(CAT)
@@ 5.2 Class-level Semantic Clustering Module
Based on the Class-level Semantic Space Hypothesis, we consider that the neighbors impact the self-attention learning of the central node with their classes as the basic units. This idea is very intuitive for heterophilic graphs; when the representations of graph nodes are difficult to distinguish, observing more nodes for a class makes it easier to obtain the global distribution of that class.
In that semantic space, we treat the local neighbors belonging to the same cluster as a whole, which is referred to as Semantic Cluster $SC=f_{\text{clustering }}(x),\quad x\in\mathbb{R}^{n\times F},\quad SC\in\mathbb{R}^{n\times 1}$. Where $SC(i)\in C$ represents the cluster class of nodes with the index $i$. Accordingly, the center of each SC in that semantic space is $SC_{c}=\left\{\frac{\sum\limits_{SC(v)=c}SC_{v_{1}}}{n_{c}},\frac{\sum\limits_{SC(v)=c}SC_{v_{2}}}{n_{c}},\ldots\frac{\sum\limits_{SC(v)=c}SC_{v_{d}}}{n_{c}}\right\},c\in C$. We can update the causal graph proposed in Figure 2 to Figure 9.
Figure 9: Causal graph behind the GAT at the Semantic Cluster level.
As shown in Figure 10, three learning paradigms can be adopted in this module to obtain Class-level Semantic Clusters. Ordered in ascending prior knowledge about node category distribution they are: unsupervised, semi-supervised, and supervised learning. The more information we acquire about the categorical distribution of graph nodes, the closer the obtained Semantic Cluster distribution will be to the distribution in the ideal semantic space. This further indicates a more accurate estimate of the total effect of attention-learning on Class-level Semantic Clusters corresponding to each category. We construct three CAT variants by adopting the following three learning paradigms in this module:
  • Unsupervised manner: For all nodes in the graph, their categorical labels are unseen. Unsupervised clustering methods can be employed to obtain a rough semantic space with the input of node features. The CAT variant built in this manner is referred to as CAT-unsup.
  • Semi-supervised manner: Categorical labels for a fixed ratio of nodes are known and used to infer the labels of unknown nodes. Classification methods with a semi-supervised setting can be employed to obtain a less rough semantic space with the input node features. The CAT variant constructed in this manner is referred to as CAT-semi.
  • Supervised manner: Categorical labels for all nodes are known and their categorical distribution is completely and accurately observed. It should be noted that the label information is only available in the Class-level Semantic Clustering stage and is not used for node classification. The CAT variant employed in this manner is referred to as CAT-sup.
@@ 5.3 Total Effect Estimation Module
As illustrated in Figure 9, there are two paths from class-level Semantic Clusters to the central node’s attention distribution that will jointly influence the representation learning of the central node. Therefore, we employed total effect as a measurement of the Distraction Effect based on the preliminary of Causal Inference. This module contains three important steps, semantic cluster intervention, intervened attention learning, and graph trimming.
@@
    Semantic Cluster Intervention. As detailed in Section 3, Total effect is estimated based on the intervention.This step is theoretically equivalent to forcing the central node to answer a causal question: how will my attention distribution change if Semantic Cluster $c$ is removed from my LND? The physical intuition behind this intervention-related question is that it is an operation that renders the nodes belonging to Semantic Cluster $c$ invisible to the central node. Figure 11 can be mathematically modelled as Eq.7, where $A$ represents the adjacency matrix of the original graph, and $\hat{A}$ represents that of the intervention graph.
@@
    Intervened Attention Learning. Since the intervention in the Semantic Cluster does not affect the shape of Class-level Semantic Space (which is governed by the data generation mechanism of the graph and considered to be invariant), we need to guarantee that before and after the intervention, the base GAT only changes its attention distribution, while the other capabilities remain unchanged. From the model implementation perspective, we do not alter the parameters responsible for transforming node features, allowing the model to solely reallocate the attention assigned to neighboring nodes and itself. The attention assigned to Semantic Clusters can be represented as Eq.8, where $\alpha_{SC}^{c}=\sum\limits_{SC(V_{j})=c}\alpha_{ij}$, and the self-attention of central node $V_{i}$ is $\alpha_{self\_attention}=1-\sum\limits_{C}\alpha_{SC}^{c}$.
@@
    Graph Trimming. According to the concept in Section 3, we can calculate the TE of Semantic Cluster $c$ based on the self-attention of the central node according to Eq.9.
@@
    The lower the value of the sum of $TE_{\alpha_{SC}}$ is, i.e., the higher the value of $TE_{\alpha_{self\_attention}}$ is, the more it can distract the central node and lead to low self-attention for the central node, and vice versa. Therefore, we remove the Semantic Cluster with lower $TE_{\alpha_{SC}}$ values and retain only the Semantic Cluster with the highest $TE_{\alpha_{SC}}$. In other words, only the Semantic Cluster with the lowest TE on self-attention of the central node will remain. Eventually, we obtain the adjacency matrix of the trimmed graph denoted as Eq.10, which equals an operation that removes the edges connecting Distraction Neighbors and the central nodes.
@@ 6.1 Databases
To ensure the richness and representativeness of the employed data, we selected seven heterophilic graphs of three sizes. All datasets possess an Edge Homophily below 0.23. The basic information of the datasets is given in the Table 2. Edge Homophily is defined in Eq.11.
@@
  • Small-size datasets. We use the WebKB dataset [33] constructed from the WebKB web page. It was collected from the computer science departments of Cornell, Texas, and Wisconsin-Madison University. This dataset was built from the hyperlinks between web pages, and the features of nodes are the bag-of-words representations. The nodes belong to five categories.
  • Medium-size datasets. We use the Chameleon and Squirrel datasets [33] collected from Wikipedia, which are applicable for node regression and node classification tasks. In these datasets, the nodes represent web pages and the edges are links between them. When applying them for the node classification tasks, the target is to predict 5 classes based on the average traffic of web pages.
  • Large-size datasets. We use the Roman-empire [43] dataset constructed from Wikipedia articles. In this dataset, the nodes represent words in a text, and the edges are constructed from their context. The target is to predict 18 classes based on the syntactic role of the nodes.
@@ 6.2 Experiments
We aim to explore the effect of neighboring nodes on the central node’s attention-learning within the GAT mechanism. Therefore, we take a GAT with fixed architecture (which can be regarded as possessing the same aggregation and feature transformation abilities) as the base model and compare its discrimination ability on graphs with different LNDs. We focus more on the difference caused by attention distribution learned by the GAT, thus choosing not to tune the architecture of the GAT through careful optimization and parameter tuning. To better demonstrate the effectiveness of CAT, we examined its efficacy across different GATs and adopted three GATs as base models. They are built with distinctive motivations and improved mechanisms and thus can present different scenarios. To distinguish between the results of different base models, we replace "G" in the original name of base models with "C" to represent their corresponding CATs. The base GATs and their fixed architecture are as follows:
  • GAT [3]. The originally proposed Graph Attention Network implicitly specifies different weights for neighboring nodes. It injects a graph structure into its self-attention strategy to learn attention coefficients, thus learning node representations in a more informative way. We set the number of GAT layers to 2 and the number of heads to 8.
  • GATv2 [44]. A dynamic graph attention variant that can learn dynamic attention by simply switching the order of internal operations in GAT. It can outperform the original GAT when more complex interactions are observed between nodes in the input graph. We set the layer of GATv2 to 2 and the number of heads to 8.
  • GATv3 [27]. A new attention mechanism that calculates the query and key from other GNN models. It can be adaptively used with homophilic or heterophilic graphs. We set the GATv3 layer to 2 and adopt a one-layer GCN in the K and Q modules. To better investigate the effect of attention-learning, we fix the weight of the calculated attention to 1 and abandon the original weighted attention strategy.
For all base GATs and their CAT variants, we use the Adam optimizer with a learning rate of 0.001 and a weight decay of 0.0001 to train the model. A single Nvidia 2080Ti GPU was used for training with a negative log likelihood loss. The maximum number of iterations was 600, and the tolerance of the early stopping strategy based on the classification accuracy on the validation set is set to 50. To evaluate the model accuracy, we divided each dataset into training, validation, and test sets at a ratio of 6:2:2 and used the average classification accuracy and standard deviation attained on the test set over 100 repetitions as the final evaluation metrics. We set the dimensions of the hidden layers to {16,32,64,128} and adopted the optimal classification accuracy. We conduct comparison and ablation experiments to verify the validity of the architecture and individual modules of CAT, respectively. Visualization experiments were also carried out to further interpret the results.
@@ 6.2.1 Comparison Experiment
We feed the original heterophilic graph and the trimmed graph obtained by CAT variants to the base GAT to obtain the final node classification accuracy. The trimmed graphs were obtained by using three variants of CAT with the following settings:
@@
  • CAT-unsup. Since the number of Semantic Clusters is known (equal to the number of target classes), we use the K-means++ algorithm in the Class-level Semantic Clustering Module in an unsupervised manner. To avoid the influence of the initial clustering centers on the results, we used 0,10,100 as random seeds for the initial clustering centers in K-means++.
@@
  • CAT-semi. For the semi-supervised manner, we employed a two-layer Multi-Layer Perception (MLP) to learn the categorical distribution of nodes. To maintain the consistency of the semi-supervised node classification task, we use the same dataset split described in Section 6.1 for the MLP.
@@
  • CAT-sup. In a supervised manner, we directly used the labels to generate the Class-level Semantic Clusters.
The results are shown in Table 3. Our approach exhibits improvements across all base GAT models. Even on the large-size dataset with an Edge Homophily level of only 0.04, the minimum relative improvements for GAT and GATv2 are 13.5% and 10.1%, respectively. Adopting semi-supervised and fully-supervised paradigms can lead to further improvements.
Additionally, we observed performance shifts on different base GATs. There are slight differences between the performances of GAT and GATv2 across most datasets. However, on the Roman-empire dataset, GATv2 outperforms GAT by over two percentage points. The potential reason could be that Roman-empire dataset contains more node categories and a larger graph size, resulting in more complex interactions between nodes in the graph, at which point the dynamic attention captured by GATv2 proves effective.
@@
GATv3 exhibits the best performance among all base models due to its incorporation of a new attention mechanism that leverages other GNN models, thereby enhancing its discrimination capability. However, CAT can further improve its classification accuracy. Among all base models, the relative improvement provided by CAT for GATv3 is the lowest. The reason is that GATv3 already boasts comparatively high discrimination capabilities on heterophilic graphs, making further enhancement more challenging. Each of the three base models exhibits strengths in different scenarios, yet CAT demonstrates the capability to further boost their performance across all datasets.
@@
In terms of the standard deviation of the prediction accuracy, on small-size datasets, the deviation of CATs is relatively large compared with that of the base GATs. However, on medium-size and large-size datasets, CAT significantly reduces the deviation and achieves more stable and statistically significant predictions.
@@
For all base GATs, CAT-sup generally outperforms CAT-unsup and CAT-semi. This is because it leverages more information in the Class-level Semantic Clustering Module, thereby obtaining a more accurate distribution of Semantic Clusters. This speculation can also explain why CAT-unsup performs worst and the CAT-semi consistently performs at a moderate level. On the one hand, this indicates that precise Class-level Semantic Clustering can facilitate better attention allocations. On the other hand, it underscores the challenge of learning better Semantic Spaces. CAT-unsup variants with different random seeds also achieve significantly different performances. For example, although all CAT-unsup models can outperform the GAT on the Wisconsin dataset, CAT-unsup with a random seed value of 10 attains an accuracy that is over 10% lower than that produced with a value of 100. CATv2-unsup exhibits a similar pattern on the Texas dataset. This indicates that we can barely guarantee that the learned Class-level Semantic Space is optimal or is approaching optimal for unsupervised learning purposes. Additionally, the results indicate that the output of Class-level Semantic Clustering plays a significant role in the overall method.
@@
Table 3: Node classification accuracy. 0, 10, and 100 represent the corresponding random seeds used in the unsupervised clustering method. The best and worst results achieved by the CAT variants are marked in bold and wavy line, respectively, and their relative improvements over GAT are shown below. The base models and the worst result among all models are marked in red. OOM represents out-of-memory.
@@ 6.2.2 Ablation Experiment
To investigate the effectiveness of each component in the proposed method, we conducted ablation studies on its two modules and accordingly obtained two trimmed graphs. For the sake of making a convincing comparison, we select CAT-unsup in this section because it performs the worst among the three variants of CAT. The results of the ablation experiment are shown in Table 4.
@@
    CAT (random_cluster). To investigate the effectiveness of the Class-level Semantic Cluster module, we replace it with a randomly assigned cluster module. We set the random seeds to 0, 10, and 100 for the random cluster assignment.
@@
    CAT(high_distraction). To investigate the effectiveness of the Total Effect Estimation, we remove the neighbors with lower distraction and create a High Distraction and Low Self-attention scenario. This model reserves the $\max\left(TE_{\alpha_{self\_attention}}\right)$ from the Total Effect Estimation.
Table 4: The results of the ablation experiments. The CAT model in this table represents CAT-unsup. The best accuracy is marked in bold, and the worst is indicated with a wavy line.
@@
CAT consistently achieves the best performance, while CAT (high_distraction) performs the worst. This comparison supports our Low distraction and High self-attention assumption and validates the efficacy of the Total Effect Estimation Module. CAT (random_cluster) gets the medium performance, indicating the significance of the Class-level Semantic Clustering Module; to a certain extent, the comparison can also aid in quantifying the impact of each class on the performance of the model. In addition, it shows that the Total Effect Estimation Module makes a larger and more stable contribution to CAT’s performance.
However, we also notice that CAT (random_cluster) can achieve results comparable to or even exceeding those of CAT in very few cases. This suggests that the clustering results obtained by the Class-level Semantic Clustering Module need optimization, whereas random clusters perform better in some instances. This phenomenon is more striking on small-size datasets, possibly because of the class imbalance issues (as shown in Figure 12), which increases the clustering difficulty.